OpenStudy (anonymous):

Find the inverse Laplace Transform F(s)= (3s+9)/(s^2+s+3)

OpenStudy (anonymous):

\[f(t)=3 e ^{-t/2}\cos[(\sqrt{11}/2)t] + (15/\sqrt{11}) e ^{-t/2} \sin[(\sqrt{11}/2)t)\] a bit tricky to solve, let me know if this is correct.